OpenStudy (anonymous):

Please help...Find the sum of the first six terms of the series: -96,144,-216,324,...

OpenStudy (mathmate):

The rule is \( a_{n+1} = -1.5 a_n \) with \( a_0=-96 , and r=-1.5 \) The partial sum is \( a_0\frac{1-r^n}{1-r}=(-96)\frac{1-(-1.5)^6}{1-(-1.5)} \)

OpenStudy (anonymous):

oohh okay thank you

OpenStudy (anonymous):

:]

OpenStudy (mathmate):

yw!